OpenStudy (anonymous):

If a circle has circumference of 81 meters, what is its diameter to the nearest meter?

OpenStudy (diyadiya):

Circumference = \(2 \times \pi \times r \)

OpenStudy (anonymous):

52 meters 118 meters 26 meters 13meters

OpenStudy (kymber):

C = 2 pi r Solve for r (radius) Plug in circumference amount Multiply by two to get diameter  :D